# Pretraining in Actor-Critic Reinforcement Learning for Robot Motion Control

### v1

@@ A.1 PPO Algorithm
We use the RSL_RL (Schwarke et al., 2025) implementation of PPO with adaptive learning rate and symmetry augmentation in RL (Mittal et al., 2024) (only for parkour tasks). The pseudo code is shown in Algorithm 1.
@@
  • When symmetry augmentation is active, new samples produced by symmetry augmentation are added to every minibatch, and the update rule is adapted accordingly. For more details, we refer the readers to Mittal et al. (2024).
@@ A.2 Dynamics Knowledge Probing Details
Pseudocode is shown in Algorithm 2. Representation extraction functions $\{f_{j}\}_{j=0}^{N-1}$ denotes the functions that extract intermediate representation from layer $j$ of one vanilla policy network, parameterized by the parameters of that policy network.
Details to note:
  • Collection by deterministic policy on state distribution tied to the stochastic policy. According to the policy parameterization formulation described in Section A.1, the parameters of policy network $\theta_{\mu}$ does not describe action uncertainty, is thus not accountable for the consequence of an action sampled stochastically from the distribution $\pi_{\theta}(o_{t})=\mathcal{N}\big(\mu_{\theta_{\mu}}(s_{t}),\sigma_{\theta_{\sigma}}\big)$. Therefore, the action-consequence pair needs to be collected by the deterministic policy $a_{t}\leftarrow\mu_{\theta_{\mu}}(o_{t})$. However, the dynamics knowledge should be assessed on the distribution of the stochastic policy $\pi_{\theta}(o_{t})$. In practice we run sufficient steps with the stochastic policy $\pi_{\theta}(o_{t})$ to make sure the states distribution fit $p_{\pi_{\theta}}(o_{t})$, and then execute mean action from deterministic policy by only one timestep and collect the transition data into dataset. We repeat this preparation-collection loop until we have enough data.
  • Considerations about lightweight MLP and limited computational budget. The correlation analysis is conducted by fitting a lightweight MLP to the action-consequence-pair dataset. If the size of the MLP and the computation is more than sufficient with respect to this task and dataset, it is probable that the networks get overfit to the training set regardless of how strong the correlation is, and thus produces extremely low error with all configuration that produces the input representation. Therefore, we intentionally use a lightweight MLP and restricted computation in all fitting runs. In practice, the trainable lightweight MLP takes the form of a 2-layer MLP, with the hidden layer’s dimension being 64 and activation function being ELU. The lightweight MLP is optimized by Adam with a learning rate of 0.001 for 20 epochs on a training set of 9000 samples.
@@
13:   $\mathcal{D}^{l}_{i}\leftarrow\{l_{t}=f_{\theta_{\mu}}(o_{t})\mid o_{t}\in\mathcal{D}_{i}\}$
@@
$\psi^{\star}=\arg\min_{\psi}\frac{1}{N}\sum_{l_{t}\in\mathcal{D}^{l}_{i}}\|g_{\psi}(l_{t})-\Delta q_{t+1}\|_{1}$
@@
$\epsilon_{ij}=\frac{1}{N}\sum_{l_{t}\in\mathcal{D}^{l}_{i}}\|g_{\psi}^{\star}(l_{t})-\Delta q_{t+1}\|_{1}$
@@ A.3 Exploration-Based Data Collection Implementation Details
where $\{\xi_{j}\}_{j=1}^{N_{I}}$ denotes $N_{I}$} individual PIDM dynamics model in the ensemble, $\hat{a_{t}}^{j}$ denotes the action inference by the $j$-th PIDM dynamics model, $\sigma_{t}$ denotes the standard deviation of predictions, $r^{i}_{t}$ indicates the intrinsic reward at timestep $t$, $c_{\text{ir}}$ is the intrinsic reward scaling factor, and $r_{\text{i\_max}}$ is the intrinsic reward clipping threshold. The two hyperparameters are tuned empirically, indicated in Table 5. The set of extrinsic reward terms are shown in Table 4.
@@
Table 4: Extrinsic reward terms in exploration-based data collection.
Item Values
Ensemble size 5
Max iterations 800
Retrain Interval 10
Retrain epochs 5
Intrinsic reward scaling factor $c_{\text{ir}}$ 10
Intrinsic reward clipping threshold $r_{\text{i\_max}}$ 30
Table 5: Exploration-based data collection hyperparameters.
@@ A.4 PIDM and Pretraining Implementation Details
Pretrain is implemented via supervised training on the pretrain dataset. Two techniques are applied to enhance robustness of the model:
@@
  • Noise addition. To increase the model’s resilience to perceptive noise, we add to every batch of training data in every epoch the noise vector sampled from the noise distribution identical to that in RL policy rollout. The scale of noise is identical across all tasks and can be found in e.g. Rudin et al. (2022). Notably, delta-state $\Delta x_{t+1}^{*}$ and ground truth action output label $a_{t}$ remain uncorrupted at all times.
@@
  • Vast transition space. With joint states, base twist, gravity, contact states, and terrain taken into consideration, the full state of a locomotion POMDP is notoriously large and thus hinders precise modeling. Recent work (Xu et al., 2025) has demonstrated success in learning a high-precision neural simulator using a lightweight GPT-2 model, which is orders of magnitude larger than our PIDM.
@@
  • Lack of inductive bias in MLP architectures might not make them the best-suited option for analysing time-series data, despite their prevalent use in RL (Bachmann et al., 2023).
Type Item Dimension
Proprioception base linear velocity 3
base angular velecity 3
projected gravity vector 3
joint position 12
joint velocity 12
[…]
Last action joint action 12
Command task specific command task-specific
Table 6: Observation space configuration
@@
PIDM-RL (Blind) Intention encoder [128, 128, 128]
@@
PIDM-RL (Perceptive) Intention encoder [512, 256, 128]
@@
Table 7: RL-PIDM architecture hyperparameters. For MLP modules, the values indicate the number of layers and number of hidder units per layer. PIDM-RL (Blind) hyperparameter set is used in blind tasks (locomotion, pedipulation), while PIDM-RL (Perceptive) hyperparameter set is used in perceptive tasks (parkour walk, crouch, jump, climb up, climb down).
A.5 RL-PIDM Implementation Details
RL-PIDM hyperparameters are also shown in Table 7.
@@ A.6 Quantitive Experiments Details
Original works are referred to as ”vanilla MLP”, and references are listed in Section 5.1. We refer the readers to their respective paper for details. Most POMDP configurations and RL hyperparameters are kept the same with the original works to the maximum possible. We especially remove some of the curriculums since they are either fixed-step (which is adapted to the original model) or make it hard to compare across runs (each curriculum step changes the reward dynamics or formulation). These are the minor modification we implement to ensure easier and fair comparison:
  • Unified collision model. The collision model of ANYmal-D of all tasks is unified as that in the main branch of Isaac Lab 2.2.
  • Pedipulation. Pedipulation now uses absolute joint action space instead of relative joint action space, because the PIDM needs the downstream tasks to operate in the same action space. Curriculum is removed, and now the environment configuration is static and identical to that of the maximum difficulty in curriculum in original work.
@@
    • Fixed-step curriculum events are removed from Climb Up and Jump, for that the curriculum steps in these 2 tasks are extremely sensitive to the timing of trigger, which disrupts the training with the new PIDM-RL architecture.
@@
    • Symmetrical augmentation (Mittal et al., 2024) is carried over from the original work. One critical observation is, adding symmetry loss does not significantly alter the training dynamics of vanilla MLP, but considerably improve the stability of PIDM-RL, especially the randomly initialized variant. So in every experiments of every method, a symmetry loss weighted by 0.2 is added. On top of this, due to the failure-prone nature of learning highly dynamic skills, some configurations are still vulnerable to producing failed runs. All configurations that did not succeed every of 5 independent runs are listed in Table 8.
@@
The choice of main performance indicator varies across tasks. Since the blind tasks (pedipulation, locomotion) are completely curriculum-free, we directly use the mean reward as the performance metric. However, since the adaptive (progress-based) terrain difficulty curriculum exist in all parkour tasks, the mean reward curve can not be directly taken as performance metric due to the mean difficulty of different methods at a certain iteration are often different. Therefore, we use the Curriculum Progress, indicated by the mean maximum terrain difficulty reached over all sub-environments as the way to assess learning performance.
More training curves that did not exist in main text are presented in Figure 8.
Figure 8: Evolution of the main performance metric during training for four of the tasks we use to evaluate. The shaded area represents the variance across five differently seeded trainings.
@@ A.7 More Ablation Experiments Training Curves
A.8 Weight Update Magnitude Observation

### v2

@@ A.1 PPO Algorithm
We use the RSL_RL (schwarke2025rslrl) implementation of PPO with adaptive learning rate and symmetry augmentation in RL (mittal2024symmetry) (only for parkour tasks). The pseudo code is shown in Algorithm 1.
@@
  • When symmetry augmentation is active, new samples produced by symmetry augmentation are added to every minibatch, and the update rule is adapted accordingly. For more details, we refer the readers to mittal2024symmetry.
@@ A.2 Dynamics Knowledge Probing Details
The experiments were performed in Pedipulation (Figure 4(b)) and Locomotion (Figure 8) tasks, each with 5 different seeds. Algorithm pseudocode is shown in Algorithm 2. Representation extraction functions $\{f_{j}\}_{j=0}^{N-1}$ denotes the functions that extract intermediate representation from layer $j$ of one vanilla policy network, parameterized by the parameters of that policy network.
Collection by deterministic policy on state distribution tied to the stochastic policy
According to the policy parameterization formulation described in Section A.1, the parameters of policy network $\theta_{\mu}$ does not describe action uncertainty, is thus not accountable for the consequence of an action sampled stochastically from the distribution $\pi_{\theta}(o_{t})=\mathcal{N}\big(\mu_{\theta_{\mu}}(s_{t}),\sigma_{\theta_{\sigma}}\big)$. Therefore, the action-consequence pair needs to be collected by the deterministic policy $a_{t}\leftarrow\mu_{\theta_{\mu}}(o_{t})$. However, the dynamics knowledge should be assessed on the distribution of the stochastic policy $\pi_{\theta}(o_{t})$. In practice we run sufficient steps with the stochastic policy $\pi_{\theta}(o_{t})$ to make sure the states distribution fit $p_{\pi_{\theta}}(o_{t})$, and then execute mean action from deterministic policy by only one timestep and collect the transition data into dataset. We repeat this preparation-collection loop until we have enough data.
Lightweight MLP formulation
The correlation analysis is performed by fitting a lightweight MLP to regress the change in joint angles from the concatenation of some internal representation and raw observations. The choice to concatenate raw observation to every intermediate representation is to ensure that, the input to the lightweight MLP contains full information about current state at all times. This makes sure that the differences between the final prediction errors are purely resulted from the understanding of dynamics of intermediate representation, not from how much it contains current state information.
Limited computation
The computation of fitting the lightweight MLP is restricted in two ways: 1) limited width and depth of MLP, and 2) limited optimizer steps. This is due to the consideration that if the size of the MLP and the computation are more than sufficient with respect to this task and dataset, the network is then able to overfit the training set regardless of how strong the correlation is, and thus produces low error with whichever input configuration. Therefore, we intentionally use a lightweight MLP and a small amount of optimizer steps in all fitting runs. In practice, the trainable lightweight MLP takes the form of a 2-layer MLP, with the only hidden layer’s dimension being 64 and activation function being ELU. The lightweight MLP is optimized by Adam with a learning rate of 0.001 for 20 epochs on a training set of 11700 samples.
Figure 8: Errors of dynamics prediction based on representations from different layers, Locomotion task.
@@
13:   $\mathcal{D}^{l}_{i}\leftarrow\{(l_{t},o_{t})\mid l_{t}=f_{\theta_{\mu}}(o_{t}),o_{t}\in\mathcal{D}_{i}\}$
@@
$\psi^{\star}=\arg\min_{\psi}\frac{1}{N}\sum_{l_{t}\in\mathcal{D}^{l}_{i}}\|g_{\psi}(l_{t},o_{t})-\Delta q_{t+1}\|_{1}$
@@
$\epsilon_{ij}=\frac{1}{N}\sum_{l_{t}\in\mathcal{D}^{l}_{i}}\|g_{\psi}^{\star}(l_{t},o_{t})-\Delta q_{t+1}\|_{1}$
@@ A.3 Exploration-Based Data Collection Implementation Details
where $\{\xi_{j}\}_{j=1}^{N_{I}}$ denotes $N_{I}$} individual PIDM dynamics model in the ensemble, $\hat{a_{t}}^{j}$ denotes the action inference by the $j$-th PIDM dynamics model, $\sigma_{t}$ denotes the standard deviation of predictions, $r^{i}_{t}$ indicates the intrinsic reward at timestep $t$, $c_{\text{ir}}$ is the intrinsic reward scaling factor, and $r_{\text{i\_max}}$ is the intrinsic reward clipping threshold. The two hyperparameters are tuned empirically, indicated in Table 4. The set of extrinsic reward terms of ANYmal D are shown in Table 5. Extrinsic reward terms of other two embodiments can be found in the supplementary code.
@@
Item Values
Ensemble size 5
Max iterations 800
Retrain Interval 10
Retrain epochs 5
Intrinsic reward scaling factor $c_{\text{ir}}$ 10
Intrinsic reward clipping threshold $r_{\text{i\_max}}$ 30
Table 4: Exploration-based data collection hyperparameters.
@@
Table 5: Extrinsic reward terms in exploration-based data collection for ANYmal-D.
@@ A.4 PIDM and Pretraining Implementation Details
Pretrain is implemented via supervised training on the pretrain dataset. In the pretraining of ANYmal D embodiment, two techniques are applied to enhance robustness of the model:
@@
  • Noise addition. To increase the model’s resilience to perceptive noise, we add to every batch of training data in every epoch the noise vector sampled from the noise distribution identical to that in RL policy rollout. The scale of noise is identical across all tasks and can be found in e.g. rudin2022learning. Notably, delta-state $\Delta x_{t+1}^{*}$ and ground truth action output label $a_{t}$ remain uncorrupted at all times.
@@
  • Vast transition space. With joint states, base twist, gravity, contact states, and terrain taken into consideration, the full state of a locomotion POMDP is notoriously large and thus hinders precise modeling. Recent work (xu2025neural) has demonstrated success in learning a high-precision neural simulator using a lightweight GPT-2 model, which is orders of magnitude larger than our PIDM.
@@
  • Lack of inductive bias in MLP architectures might not make them the best-suited option for analysing time-series data, despite their prevalent use in RL (bachmann2023scaling).
@@
PIDM (RL-Blind) Intention encoder [128, 128, 128]
@@
PIDM (RL-Perceptive) Intention encoder [512, 256, 128]
@@
Table 6: PIDM architecture hyperparameters. The values enclosed in square brackets indicate the number of layers and number of hidder units per layer in the corresponding MLP modules. PIDM (RL-Blind) hyperparameter set is used in blind tasks (locomotion, pedipulation), while PIDM (RL-Perceptive) hyperparameter set is used in perceptive tasks (parkour walk, crouch, jump, climb up, climb down).
A.5 PIDM Implementation Details
PIDM hyperparameters are also shown in Table 6.
@@ A.6 Quantitive Experiments Details
Group Members Tasks Models
Blind Perceptive Vanilla MLP PIDM
Proprioception base linear velocity
single
timestep
multiple
timesteps
base angular velocity
projected gravity vector
joint position
joint velocity
Exteroception height scan - -
Last action joint action
single
timestep
multiple
timesteps
Command task specific command
Table 7: Observation space configuration. “Blind” tasks: locomotion, pedipulation. “Perceptive” tasks: parkour walk, climb up, climb down, crouch, and jump.
Original works are referred to as “vanilla MLP”, and references are listed in Section 5.1. Most POMDP configurations and RL hyperparameters are kept the same with the original works to the maximum possible, and details can be found in the corresponding references.
A.6.1 Summary of Observation Space, Action Space and Simulation
Observation space.
All possible components of observation space are listed in Table 7. To summarize, there are 2 types of variation in the composition of observation space across all the runs: a) for the blind tasks (locomotion, pedipulation), policies do not have access to exteroception (height scans), while exteroception is included in the rest perceptive tasks. b) Vanilla MLPs do not have access to history proprioceptions as in the original works, while the inputs to PIDM based models contain history proprioceptions. The necessity of including a history of proprioception for PIDM is mainly due to the absense of terrain information and contact state in proprioception, and the presence of noise and domain randomization techniques in the training process (during both exploration-based data collection and training of task-specific policies). Therefore, it would be inappropriate to fit the PIDM with only one single frame of current proprioceptive state, due to the fact that one certain proprioception can be observed in a range of actual full states in the POMDP. In addition, there is a line of works (ji2022concurrent; portela2025whole) that suggest that proprioception history is informative and a number of useful values (e.g. foot height, contact probability, end effector force) can be estimated from it.
Action space.
The action space is the target joint position (relative to default joint positions) command that will be sent to actuator-nets (hwangbo_learning_2019) of ANYDrive 4.0.
Simulation.
All training is performed in Isaac Lab (mittal2025isaac) using 4,096 parallel environments, each running 24 simulation steps per RL iteration, for a total of 98,304 environment steps per iteration. Each environmental step corresponds to 5 ms in real time and is computed using 4 physics solver steps.
A.6.2 Modifications
Following modifications are made to ensure easier and fair comparison.
Unified collision model.
The collision model of ANYmal-D of all tasks is unified as that in the main branch of Isaac Lab 2.2.
Pedipulation.
We change the action space of Pedipulation task from relative joint action space to absolute joint action space, to make it consistent with all other tasks. The curriculum (gradual expansion of the space where command is sampled) is removed because it makes comparing rewards from different stages/runs not appropriate. After removal of curriculum, the environment configuration is static and identical to that of the maximum difficulty in curriculum in original work.
@@ A.6.2 Modifications
  • Fixed-step curriculum events are removed from Climb Up and Jump, for that the curriculum steps in these 2 tasks are extremely sensitive to the timing of trigger, which interferes with the learning once the architecture of networks changes.
@@
  • Symmetrical augmentation (mittal2024symmetry) is carried over from the original work. One critical observation is, adding symmetry loss does not significantly alter the training dynamics of vanilla MLP, but considerably improve the stability of RL training of PIDM, especially the randomly initialized variant. So in every experiments of every method, a symmetry loss weighted by 0.2 is added. On top of this, due to the failure-prone nature of learning highly dynamic skills, some configurations are still vulnerable to producing failed runs. All configurations that did not succeed every of 5 independent runs are listed in Table 8.
@@
A.6.3 Performance Indicator
The choice of main performance indicator varies across tasks. Since the blind tasks (pedipulation, locomotion) are completely curriculum-free, we directly use the mean reward as the performance metric. However, since the adaptive (progress-based) terrain difficulty curriculum exist in all parkour tasks, the mean reward curves can not be directly taken as performance metric of policies because the evolving terrain difficulty. Therefore, we use the Curriculum Progress, indicated by the average of maximum terrain difficulty reached over all sub-environments as the way to assess learning performance.
A.6.4 More Reinforcement Learning Training Curves Corresponding to Presented Quantitative Results
See Figure 9.
Figure 9: Evolution of the main performance metric during training for Parkour Walk, Pedipulation, Crouch and Jump tasks with ANYmal D. The shaded areas denote standard deviations across five seeds.
Figure 10: Evolution of the mean reward during learning velocity-tracking locomotion with Unitree Go1 (quadruped) and Unitree G1 (humanoid). The shaded areas denote standard deviations across five seeds.
@@ A.7 More Ablation Experiments Training Curves
A.8 Study of Relation between Error Level of PIDM and RL Performance
(a) 75%
(b) 60%
(c) 40%
Figure 13: PIDM error levels of the 3 checkpoints used to study the relation between PIDM error and RL performance, indicated by absolute and normalized joint errors.
Figure 14: Study of relation between error Level of PIDM and RL performance.
To verify the positive correlation between the accuracy of pretrained PIDM and resulted gain in RL performance, we study 3 PIDM model checkpoints from pretraining which respectively produces normalized error level of 75%, 60%, and 40%, as shown in Figure 13. The checkpoint that produces normalized error of 40% is the one that used to present the main quantitative results.
We benchmarked the 3 model checkpoints, along with the randomly initialized PIDM in the Climb Down and Climb Up tasks and plotted the performance indicator curves in Figure 14. The results suggest that in both tasks PIDM of 40% performs the best among all variant, followed by the model of 60% error. PIDM of 75% and randomly initialized PIDM are relatively underperforming the other two. These results suggest that the more accurate PIDM tends to yield better performance in downstream RL tasks. This finding might also hint that larger benefits are possible if higher accuracy can be attained using larger model and more modern architectures to implement PIDM. We leave this for future work.
A.9 Weight Update Magnitude Observation
@@
A.10 Sim-to-Real Transfer Demonstration
We have deployed the policy Locomotion on a real ANYmal D robot. The video is attached in the supplementary material.